OpenStudy (anonymous):

plz need help fast(will give medal+fan) A volleyball playoff competition has eliminations each round. The table below shows the number of teams in each round of the tournament: Round (x) 1 2 3 4 Teams f(x) 16 8 4 2 Compute the average rate of change of f(x) from x = 1 to x = 4 and identify the meaning of that rate. negative three over fourteen; on average, there were 3 fewer teams every 14 rounds negative fourteen over three; on average, there were 14 fewer teams every 3 rounds 3; there were 3 rounds between rounds 1 and 4 14; there were 14 fewer team between rounds 1 and 4

OpenStudy (anonymous):

@BabyGirlAj can u help

OpenStudy (anonymous):

I will try

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Round (x) 1 2 3 4 Teams f(x) 16 8 4 2 [f(4) - f(1)] / (4 - 1 ) = ( 2 - 16) / ( 3) = -14/3 negative fourteen over three; on average, there were 14 fewer teams every 3 rounds

OpenStudy (anonymous):

thank u much : )

OpenStudy (anonymous):

Anytime :)